OpenStudy (anonymous):

I need help with Applied Calculus 1, Find dy/dx by implicit differentiation. (x^2)-6xy+(y^2)=6

OpenStudy (anonymous):

\[x^2-6xy+y^2=6\]

OpenStudy (freckles):

hey

OpenStudy (anonymous):

hello

OpenStudy (freckles):

so can i see you try for differentiating both sides w.r.t. x?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

\[2x-6(y+x \frac{ dy }{ dx })+2y(\frac{ dy }{ dx })=0\] \[2x-6y-6x \frac{ dy }{ dx }+2y \frac{ dy }{ dx }\] \[-6x \frac{ dy }{ dx }+2y \frac{ dy }{ dx }=6y+2x\] \[\frac{ dy }{ dx }=\frac{ 6y+2x }{ 2y-6x }\]

OpenStudy (freckles):

\[\frac{d}{dx}x^2=2x \\ \frac{d}{dx}(-6xy)=-6 \frac{d}{dx}(xy)=-6(y \frac{dx}{dx} +x \frac{dy}{dx}) =-6(y+x \frac{dy}{dx}) \\ \frac{d}{dx}y^2=2y \frac{dy}{dx}\] the differentiating part looks great checking your algebra...

OpenStudy (freckles):

only complaint is that when you subtract 2x on both sides you put +2x on the right hand side

OpenStudy (anonymous):

ooh your right

OpenStudy (anonymous):

totally caught me off guard

OpenStudy (anonymous):

so everything looks good except for the sign right?

OpenStudy (freckles):

yep

OpenStudy (anonymous):

awesome thanks!

OpenStudy (freckles):

well

OpenStudy (freckles):

you could do a little simplifying if you wanted

OpenStudy (freckles):

but yeah the answer you have is right

OpenStudy (anonymous):

3y-x/y-3x?

OpenStudy (freckles):

with the sign change

OpenStudy (freckles):

\[\frac{dy}{dx}=\frac{6y-2x}{2y-6x} =\frac{2(3y-x)}{2(y-3x)}=\frac{3y-x}{y-3x}\]

OpenStudy (anonymous):

yup

OpenStudy (freckles):

but that is all you can really do